OpenStudy (anonymous):

If f(1)=12, f'(x) is continuous and the integral from 1 to 4 of f'(x)dx=17, what is the value of f(4)?

OpenStudy (anonymous):

If \[f(1)=12\] f' is continuous and \[\int\limits_{1}^{4}f'(x)dx=17\] what is the value of f(4)?

OpenStudy (shubhamsrg):

well,,since f' is continuous, you wouldnt have to break the limits of the integral anywhere.. and integral(f'x)dx from 1 to 4 is obviously f(4) - f(1) and its given f(1)= 12 f(4)-f(1)=17 just solve for f(4)